UT 20120405 AO#2 Commissioning

Observer: ARiccardi, LBusoni, FQuiros-Pacheco, CArcidiacono (LBTO)
Telescope Operator: GBechetti (LBTO)
Additional AO Support: DMiller
Telescope Support: MMidkiff(LBTO)
Instrument Support: DOfficer (LBTO)

Summary

Details

03:00 Dome open, closed AO Loop on a bright star

03:20 Try to set pointing base in PCS so Target star land on the center of the narrow field of view. Troubles with closed loop offsets.

04:04 Spent 1.3 hours trying to get offsets in AO Closed loop to work. Talked to Chris B. on the phone, but cause not found. IT#4023

04:40 Continue without offset via IIF. Lorenzo perform offset from AO Engineering GUI. Collect data with 500 modes needed to characterize the ND filter

05:15 Move to Zenith and insert ND in front of AO system. Restart PCS

05:30 Closed loop with bin 2 to characterize ND.

-- DougMiller - 05 Apr 2012
Topic revision: r1 - 05 Apr 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback